# LLM-3D Print: Large Language Models To Monitor and Control 3D Printing

### v2

@@ Abstract
Industry 4.0 has revolutionized manufacturing by driving digitalization and shifting the paradigm toward additive manufacturing (AM). Fused Deposition Modeling (FDM), a key AM technology, enables the creation of highly customized, cost-effective products with minimal material waste through layer-by-layer extrusion, posing a significant challenge to traditional subtractive methods. However, the susceptibility of material extrusion techniques to errors often requires expert intervention to detect and mitigate defects that can severely compromise product quality. While automated error detection and machine learning models exist, their generalizability across diverse 3D printer setups, firmware, and sensors is limited, and deep learning methods require extensive labeled datasets, hindering scalability and adaptability. To address these challenges, we present a process monitoring and control framework that leverages Large Language Models (LLMs) alongside 3D printers to detect and address printing defects—entirely without domain-specific fine-tuning or training. Instead, the LLM operates using in-context learning and general reasoning to interpret print quality from images captured after each layer or print segment, identify failure modes, and query the printer for relevant parameters. It then autonomously generates and executes a corrective action plan. We validated the effectiveness of the proposed framework in identifying defects by comparing it against a control group of engineers with diverse AM expertise. Our evaluation demonstrated that LLM-based agents not only accurately identify common 3D printing errors—such as inconsistent extrusion, stringing, warping, and layer adhesion—but also effectively determine the underlying causes and autonomously correct them without any need for human intervention.
@@ 1 Introduction
Industry 4.0, also known as the Fourth Industrial Revolution, integrates the Internet of Things (IoT), artificial intelligence, and big data analytics to assist in smart manufacturing where machines and systems seamlessly communicate to optimize production processes 1, 2, 3. This technological integration enhances quality control and certification processes via in-situ process monitoring utilizing automated data collection to ensure consistent adherence to quality standards 4. Automated documentation facilitates comprehensive traceability, simplifies certification, and provides verifiable records that help manufacturers achieve and maintain industry certifications, boosting compliance and customer trust 5.
@@
Additive Manufacturing (AM) plays a key role in this industrial revolution, enabling rapid prototyping from conceptual design to production level parts through an iterative design process bypassing the constraints of conventional manufacturing processes 6. AM significantly reduces production time, enhances design flexibility, lowers costs, minimizes post-processing, and supports multi-material design through a layer-by-layer printing process7. This gives AM a substantial advantage over traditional subtractive manufacturing techniques, such as CNC machining, by offering greater material efficiency and design versatility 8. Consequently, AM has shown promising applications in various fields, including healthcare 9, 10, medical devices 11, 12, and aerospace 13, among others 14, 15, 16.
@@
One of the most widely used additive manufacturing processes is Fused Deposition Modeling (FDM). This technique involves extruding thermoplastic filaments through a heated nozzle, which deposits material layer by layer to construct a part from the bottom up 17. The accessibility and versatility of FDM have fueled the growth of open-source projects like RepRap 18, 19, democratizing 3D printing technology and enabling individuals, small businesses 20, and research labs 21, 22 to develop and customize hardware, significantly expanding the applications and accessibility of additive manufacturing.
@@
Despite its advantages, FDM is highly susceptible to errors and failures, which limit its consistency and material efficiency. Common issues include warpage, layer misalignment, under-extrusion, over-extrusion, and stringing, among others. 23, 24, 25, 26, 27 These issues are further compounded by the lack of standardized hardware, material variations 28, user errors 29, and print process parameters that are not tailored to specific designs 30, 31, all of which contribute to significant variations in print quality. Studies indicate that 20% of PLA prints fail 32, ABS prints have a material waste rate of 34% 33, and overall failure rates can reach 41.1%, with human error accounting for 26.3% of these failures 29. These failures not only result in wasted material, energy, and time but also limit the use of AM parts in end-use products, especially in safety-critical applications like medical devices and aerospace components.
@@
Some approaches address print failures by simulating the 3D printing process and optimizing part-specific parameters 34, 30. While effective, these simulations are computationally expensive and impractical for production, especially when compared to the lower inference cost/time of LLMs, which operate on a per-layer basis in seconds rather than hours. Another strategy targets the limitations of open-loop FDM systems by introducing closed-loop control, where “real-time” refers to parameter adjustments made at the end of each print layer or even a more finer sampling rate. Prior work has incorporated sensors such as accelerometers 35, acoustic sensors 36, and laser scanners 37 to detect anomalies. However, many defects, such as stringing or subtle layer separation, lack strong physical signatures and remain difficult to detect with such sensors. Additionally, the cost and complexity of these sensor systems hinder their widespread use 38, motivating more generalizable, multimodal approaches like the one proposed here.
@@
Recent research has leveraged machine learning for error detection in FDM, demonstrating promising results through camera-based monitoring and convolutional neural networks to predict and correct issues such as extrusion rate 58, warpage 59, surface defects 60, and layer defects 61. One notable approach aims to generalize the process across different parts, materials, and printing systems by using convolutional neural networks to detect errors and adjust parameters like flow rate, speed, z-offset, and extruder temperature 62.
However, most methods are limited to addressing errors in a single modality by changing only a few print parameters and have primarily demonstrated correction of the flow rate parameter in a single geometry used for both training and testing. Additionally, some machine learning models require a reference object for comparison, which restricts their effectiveness for custom parts 60. Furthermore, these methods are not capable of real-time correction, meaning that if an error is detected, the part cannot be recovered. Although Brion et al. 62 addressed this by splitting the toolpath for a layer into smaller segments, optimizing one segment does not necessarily ensure that subsequent segments will also be optimized, especially if their shapes differ significantly.
@@
Leveraging transformer architecture 63, 64 and massive datasets 65, 66, 67, 68, 69, 70, Large Language Models (LLMs) have made significant strides in various natural language processing (NLP) tasks, including text generation and following task-specific instructions 71, 72. Additionally, LLMs have demonstrated emergent reasoning capabilities by interpolating and utilizing their extensive training data, allowing them to make inferences, draw conclusions, and solve problems beyond their explicit programming 73, 74. Despite the substantial computational resources required for training and fine-tuning LLMs for specific applications, these models have shown an exceptional ability to generalize to new tasks and domains. This remarkable capability is primarily attributed to the in-context learning (ICL) paradigm 65. By using minimal natural language prompts and avoiding extensive fine-tuning, LLMs have emerged as effective ”few-shot learners” 75, 76, demonstrating proficiency with limited training examples.
@@
The adaptability of Large Language Models (LLMs) to new domains and their ability to learn in context has been transformative across numerous scientific fields. In chemistry, for example, LLMs have autonomously designed, planned, and executed complex experiments 77, 78. In mathematics and computer science, these models have discovered novel solutions to longstanding problems like the cap set problem and optimized algorithms for challenges such as the bin-packing problem 79. Furthermore, LLMs have significantly advanced research in biomedical fields 80, 81, 82, materials science 83, 84, and environmental science 85. Their impact extends to other scientific domains as well, enhancing understanding and expanding capabilities 86, 87, 88, 89, 90. Additionally, LLMs have proven to be effective optimizers for foundational problems such as linear regression and the traveling salesman problem, often matching or exceeding the performance of specialized heuristics through straightforward prompting 91. This versatility and efficacy underscore the potential of LLMs to drive innovation and efficiency in a wide array of scientific and engineering applications.
@@
In the field of mechanical engineering, fine-tuned Large Language Models (LLMs) have demonstrated exceptional capabilities in advanced tasks such as knowledge retrieval, hypothesis generation, and agent-based modeling. They have played a crucial role in integrating diverse domains through the use of knowledge graphs 92, 93. Additionally, LLMs excel in various design-related tasks, including sketch similarity analysis, material selection, engineering drawing analysis, CAD generation, and structural optimization 94, 95. Their ability to seamlessly connect and enhance these tasks highlights their transformative potential in mechanical engineering.
@@
Figure 2: Schematic of the Proposed Framework. The process begins with a G-code file (a) being uploaded to the 3D printer, (b) which is equipped with two frame-mounted cameras. After each layer is printed, the extruder moves to the home position, and two images of the current print state are captured (c). These images are analyzed by the LLM, which evaluates the print, makes observations, and identifies any failures. If failures are detected, the LLM supervisor (d) invokes the information planner (e). The executor (f) then carries out the information gathering plan, after which the solution planner (g) is activated by the supervisor. The solution plan is executed by another executor (h), and finally, the supervisor invokes the handoff module (i) to resume the print.
@@
The hierarchical machine-to-machine framework operates by capturing two images of the ongoing 3D print—one from the top and the other from the front once a layer is completed and the print is paused. These images, along with the description of the part, are fed into an LLM, which evaluates the print quality, identifies defects, and makes relevant observations.
Upon identifying an issue, the supervisory LLM invokes a planner to generate a detailed plan outlining the necessary information and queries for the printer to diagnose the problem. Another LLM agent then executes tools to retrieve this information via the printer’s API. Based on the gathered data, the LLM generates a solution plan, which is implemented by another agent through direct communication with the printer’s API. After the solution is executed, the supervisory LLM verifies the parameters and resumes the print. The key advantage of this supervisory LLM is its ability to track the entire conversation among all LLM agents and orchestrate their actions as needed. This comprehensive oversight ensures system coordination and efficiency, with each agent being invoked precisely when necessary.
One of the significant benefits of this framework is its flexibility to work across various 3D printers, optimizing print parameters specific to each part without requiring any pre-existing dataset. This adaptability allows the system to fine-tune process parameters on-the-fly, accommodating different materials, geometries, and printer settings. Additionally, the LLM provides detailed process commentary, aiding in the certification of the part and enhancing traceability. This increases trust in the final product by comprehensively documenting the manufacturing process, reducing the need for destructive testing to validate part integrity. Real-time, insightful commentary on the printing process supports quality assurance and ensures compliance with industry standards and regulations.
@@ 2 Methodology
Figure 3: Schematic of the Proposed Framework. The process begins with a G-code file (a) being uploaded to the 3D printer, which is equipped with two frame-mounted cameras (b). After each layer is printed, the extruder moves to the home position, and two images of the current print state are captured (c). These images are analyzed by the LLM, which evaluates the print, makes observations, and identifies any failures. If failures are detected, the LLM supervisor (d) invokes the information planner (e). The executor (f) then carries out the information gathering plan, after which the solution planner (g) is activated by the supervisor. The solution plan is executed by another executor (h), and finally, the supervisor invokes the handoff module (j) to resume the print.
Given the complexity of the task, our multi-agent LLM framework consists of seven key modules, each designed to perform specific tasks independently as illustrated in Figure 2. The framework includes an Image-Based Reasoning Module (Figure 2c) that identifies defects and makes observations when the printer is paused. There are two planning modules, one focusing on gathering necessary information (Figure 2e) and the other on developing solutions to the identified issues (Figure 2g). Additionally, three execution modules interact directly with the 3D printer to apply these solutions (Figure 2f, h, i). A Supervisor Module (Figure 2d) oversees the entire operation, maintaining a dynamic state dictionary that is accessible to other modules, facilitating agent communications, and orchestrating the timely activation of modules based on the current data and needs. By leveraging the specialized abilities and collaborative dynamics among these agents, our multi-agent system not only enhances efficiency in tackling complex tasks but also supports the integration of additional functionalities, allowing for scalable and adaptable enhancements as needed.
@@ 2.1 Error Detection
Large Vision-Language Models (LVLMs) such as MiniGPT-4 96 and LLaVA 97 have demonstrated significant capabilities in understanding and interpreting images, achieving remarkable performance across various visual tasks 98, 99, 100, 101. Owing to extensive pre-training on large datasets of image-text pairs, these models possess exceptional generalization abilities, enabling impressive zero-shot performance. They can effectively recognize and process both familiar and unfamiliar objects without the need for additional training data. This capability to generalize from diverse datasets makes them highly versatile and powerful tools for a wide range of visual and multimodal applications.
@@
Our methodology employs a finely tuned prompt (SI 1) that instructs the LLM to act as a 3D printing expert agent. The LLM evaluates print quality, identifies visible failure modes, and provides structured, focused, and actionable responses in specified format.

### v3

@@ Abstract
Industry 4.0 has revolutionized manufacturing by driving digitization and shifting the paradigm toward additive manufacturing (AM). Material extrusion (MEX), a core AM method, produces customized and cost-effective products with minimal waste, challenging traditional subtractive manufacturing. Despite its advantages, MEX remains susceptible to defects that can compromise part quality and function, often requiring expert intervention. Existing rule-based and machine learning approaches struggle to generalize across different printers and sensors, while deep learning methods depend on large labeled datasets, limiting their scalability and adaptability. To address these challenges, we introduce a process monitoring and control framework that employs Large Language Models (LLMs) as autonomous controllers for additive manufacturing. Unlike rule-based or heavily data-dependent approaches, our method requires no domain-specific fine-tuning or training. Instead, the LLM leverages in-context learning, self-prompting, and iterative prompt-reason refinement to evaluate print quality from sequential image captures, detect and classify emerging failure modes, and query and modify the printer for relevant operating parameters. Through this adaptive reasoning process, the LLM not only interprets defects but also improves its own decision-making logic, autonomously formulating and executing corrective actions. This demonstrates a rule-free, self-improving approach to process control that extends beyond traditional quality assurance methods. We validated the effectiveness of the proposed framework by comparing it with a control group of engineers with different levels of AM expertise. The evaluation showed that LLM-based agents not only reliably identified common 3D printing errors such as inconsistent extrusion, stringing, warping, and poor layer adhesion, but also determined their causes and corrected them without human intervention. In addition to matching expert-level accuracy, the LLM was able to recognize emerging print errors earlier than human experts, highlighting its value as a proactive controller. To further demonstrate generalizability, we deployed and tested the framework on two different 3D printers with distinct sensor setups, confirming its adaptability across hardware. We also performed compression tests on baseline prints and on prints optimized by the LLM, with the optimized parts showing clear improvements in mechanical performance.
@@ 1 Introduction
Industry 4.0, also known as the Fourth Industrial Revolution, integrates the Internet of Things (IoT), artificial intelligence, and big data analytics to assist in smart manufacturing where machines and systems seamlessly communicate to optimize production processes 1, 2. This technological integration enhances quality control and certification processes via in-situ process monitoring, utilizing automated data collection to ensure consistent adherence to quality standards 3. Automated documentation facilitates comprehensive traceability, simplifies certification, and provides verifiable records that help manufacturers achieve and maintain industry certifications, boosting compliance and customer trust 4.
@@
Additive Manufacturing (AM) plays a key role in this industrial revolution, enabling rapid prototyping from conceptual design to production level parts through an iterative design process, bypassing the constraints of conventional manufacturing processes 5. AM significantly reduces production time, enhances design flexibility, lowers costs, minimizes post-processing, and supports multi-material design through a layer-by-layer printing process6. This gives AM a substantial advantage over traditional subtractive manufacturing techniques, such as CNC machining, by offering greater material efficiency and design versatility 7. Consequently, AM has shown promising applications in various fields, including healthcare 8, 9, medical devices 10, 11, and aerospace 12, among others 13, 14, 15.
@@
One of the most widely used additive manufacturing processes is Material Extrusion (MEX). This technique involves extruding thermoplastic filaments through a heated nozzle, which deposits material layer by layer to construct a part from the bottom up 16. The accessibility and versatility of MEX have fueled the growth of open-source projects like RepRap 17, 18 and Klipper19. These community driven projects have democratized 3D printing and have enabled individuals, small businesses 20, and research labs 21, 22 to develop and customize hardware, significantly expanding the applications and accessibility of additive manufacturing.
@@
Despite its advantages, MEX is highly susceptible to errors and failures, which limit its consistency and material efficiency. Common issues include warpage, layer misalignment, under-extrusion, over-extrusion, and stringing, among others. 23, 24, 25, 26, 27 These issues are further compounded by the lack of standardized hardware, material variations 28, user errors 29, and print process parameters that are not tailored to specific designs 30, 31, all of which contribute to significant variations in print quality. Studies indicate that 20% of PLA prints fail 32, ABS prints have a material waste rate of 34% 33, and overall failure rates can reach 41.1%, with human error accounting for 26.3% of these failures 29. These failures not only result in wasted material, energy, and time but also limit the use of AM parts in end-use products, especially in safety-critical applications like medical devices and aerospace components.
@@
Some approaches address print failures by simulating the 3D printing process and optimizing part-specific parameters 34, 30. While effective, these simulations are computationally expensive and impractical for production, especially when compared to the lower inference cost/time of LLMs, which operate on a per-layer basis in seconds rather than hours. Another strategy targets the limitations of open-loop MEX systems by introducing closed-loop control, where “real-time” refers to parameter adjustments made at the end of each print layer or even a more finer sampling rate. Prior work has incorporated sensors such as accelerometers 35, acoustic sensors 36, and laser scanners 37 to detect anomalies. However, many defects, such as stringing or subtle layer separation, lack strong physical signatures and remain difficult to detect with such sensors. Additionally, the cost and complexity of these sensor systems hinder their widespread use 38, motivating more generalizable, multimodal approaches.
@@
Recent research has leveraged machine learning for error detection in MEX, demonstrating promising results through camera-based monitoring and convolutional neural networks to predict and correct issues such as extrusion rate 58, warpage 59, surface defects 60, and layer defects 61. One notable approach aims to generalize the process across different parts, materials, and printing systems by using convolutional neural networks to detect errors and adjust parameters such as flow rate, speed, z-offset, and extruder temperature 62.
However, most of these methods are limited to addressing errors within a single modality by changing only a few print parameters and have primarily demonstrated correction of the flow rate parameter inside a single geometry used in both training and testing. Additionally, some machine learning models require a reference object for comparison, which restricts their effectiveness for custom parts 60. Furthermore, these methods are not capable of real-time correction, meaning that if an error is detected, the part cannot be recovered. Although Brion et al. 62 addressed this by splitting the toolpath for a layer into smaller segments, optimizing one segment does not necessarily ensure that subsequent segments will also be optimized, especially if their shapes differ significantly.
@@
Leveraging transformer architecture 63, 64 and massive datasets 65, 66, 67, 68, 69, 70, LLMs have made significant strides in various natural language processing (NLP) tasks, including text generation and following task-specific instructions 71, 72. Additionally, LLMs have demonstrated emergent reasoning capabilities by interpolating and utilizing their extensive training data, allowing them to make inferences, draw conclusions, and solve problems beyond their explicit programming 73, 74. Despite the substantial computational resources required for training and fine-tuning LLMs for specific applications, these models have shown an exceptional ability to generalize to new tasks and domains. This remarkable capability is primarily attributed to the in-context learning (ICL) paradigm 65. By using minimal natural language prompts and avoiding extensive fine-tuning, LLMs have emerged as effective ”few-shot learners” 75, 76, demonstrating proficiency with limited training examples.
@@
The adaptability of LLMs to new domains and their ability to learn in context has been transformative across numerous scientific fields. In chemistry, for example, LLMs have autonomously designed, planned, and executed complex experiments 77, 78. In mathematics and computer science, these models have discovered novel solutions to longstanding problems such as the cap set problem and optimized algorithms for challenges such as the bin-packing problem 79. Furthermore, LLMs have significantly advanced research in biomedical fields 80, 81, 82, materials science 83, 84, and environmental science 85. Their impact extends to other scientific domains as well, enhancing understanding and expanding capabilities 86, 87, 88, 89, 90. Additionally, LLMs have proven to be effective optimizers for foundational problems such as linear regression and the traveling salesman problem, often matching or exceeding the performance of specialized heuristics through straightforward prompting 91. This versatility and efficacy underscore the potential of LLMs to drive innovation and efficiency in a wide array of scientific and engineering applications.
@@
In the field of mechanical engineering, fine-tuned LLMs have demonstrated exceptional capabilities in advanced tasks such as knowledge retrieval, hypothesis generation, and agent-based modeling. They have played a crucial role in integrating diverse domains through the use of knowledge graphs 92, 93. Additionally, LLMs excel in various design-related tasks, including sketch similarity analysis, material selection, engineering drawing analysis, CAD generation, and structural optimization 94, 95. Their ability to seamlessly connect and enhance these tasks highlights their transformative potential in mechanical engineering. Multi-Modal Vision-language models (VLMs), which combine visual and textual inputs, are now extending these capabilities by supporting parametric CAD sketch generation via CadVLM, which uses sketch images and primitive/text constraints to perform autocompletion, autoconstraint, and conditional image-based CAD sketch generation with high accuracy 96. In manufacturing contexts VLMs are also being used to recognize manufacturing features in CAD models without relying heavily on predetermined geometric rules, by using techniques like few-shot learning, multi-view prompts, and sequential reasoning to generalize across processes such as machining, additive manufacturing, sheet metal forming, mold-making, and casting 97.
In additive manufacturing specifically, LLMs and VLMs are fundamentally transforming the 3D printing ecosystem across multiple critical domains. Comprehensive evaluations reveal that VLMs exhibit exceptional versatility throughout the complete engineering design pipeline, from initial conceptualization to final manufacturing and quality inspection, with particularly notable efficacy in design for manufacturing (DfM) evaluations tailored to additive processes 98. This has led to the development of specialized manufacturing-oriented VLMs such as MaViLa, which leverages retrieval-augmented generation (RAG) architectures to integrate domain-specific manufacturing knowledge, achieving demonstrably superior performance compared to general-purpose VLMs across critical manufacturing applications 99.
@@
Figure 2: Schematic of the Proposed Framework. (a) The process begins with uploading a part’s G-code file to the 3D printer. (b) The printer is equipped with two frame-mounted cameras. (c) After each layer is printed, the extruder moves to the home position, and two images of the current print state are captured. These images are analyzed by the LLM, which evaluates the print, makes observations, and identifies any failures. (d, e) If failures are detected, the LLM supervisor invokes the information planner. (f) The executor then carries out the information gathering plan. (g) Afterwards, the solution planner is activated by the supervisor. (h) The solution plan is executed by another executor. (i, j) Loop ends with the supervisor invoking the handoff module to resume the print.
@@
The hierarchical machine-to-machine framework operates by capturing two images of the ongoing 3D print, one from the top and the other from the front, once a layer is completed and the print is paused. These images, along with the natural language description of the part, are fed into an LLM, which evaluates the print quality, identifies defects, and makes relevant observations.
Upon identifying an issue, the supervisory LLM invokes a planner to generate a detailed plan outlining the necessary information and queries for the printer to diagnose the problem. Another LLM agent then executes tools to retrieve this information via the printer’s API. Based on the gathered data, the LLM generates a solution plan, which is implemented by another agent through direct communication with the printer’s API. After the solution is executed, the supervisory LLM verifies the parameters and resumes the print. The key advantage of this supervisory LLM is its ability to track the entire conversation among all LLM agents and orchestrate their actions as needed. This comprehensive oversight ensures system coordination and efficiency, with each agent invoked precisely when necessary.
One of the significant benefits of this framework is its flexibility to work across various 3D printers, optimizing print parameters specific to each part without requiring any pre-existing dataset. This adaptability allows the system to fine-tune process parameters dynamically, accommodating different materials, geometries, and printer settings. Additionally, the LLM provides detailed process commentary, aiding in the certification of the part and enhancing traceability. This increases trust in the final product by comprehensively documenting the manufacturing process, reducing the need for destructive testing to validate part integrity. Real-time, insightful commentary on the printing process supports quality assurance and ensures compliance with industry standards and regulations.
@@ 2 Methodology
Figure 3: Prompt Pipeline for individual agents. (a) Error Detection Agent – Images from the paused printer (1) are formatted into a (2) predefined prompt and passed to the (3) VLM. The VLM analyzes them to assess print quality and detect failures, returning (4) a structured response. (b) Planning Agent – An adapter module uses an LLM to refine and adapt reasoning prompts based on outputs from the Error Detection Agent. These prompts, together with existing information, guide the planner in formulating an actionable plan. (c) Executor Agent – The LLM selects the appropriate API endpoint to execute the plan on the printer. It validates the command’s success against the printer’s response, adjusting as needed. Internal reasoning and tool use are employed throughout this stage.
Given the complexity of the task, our multi-agent LLM framework consists of seven key modules, each designed to perform specific tasks independently as illustrated in Fig.2. The framework includes an Image-Based Reasoning Module (Fig.2c) that identifies defects and makes observations when the printer is paused. There are two planning modules, one focusing on gathering necessary information (Fig.2e) and the other on developing solutions to the identified issues (Fig.2g). Additionally, three execution modules interact directly with the 3D printer to apply these solutions (Fig.2f, 2h, 2i). A Supervisor Module (Fig.2d) oversees the entire operation, maintaining a dynamic state dictionary that is accessible to other modules, facilitating agent communications, and orchestrating the timely activation of modules based on the current data and needs. By leveraging the specialized abilities and collaborative dynamics among these agents, our multi-agent system not only enhances efficiency in tackling complex tasks but also supports the integration of additional functionalities, allowing for scalable and adaptable enhancements as needed.
@@ 2.1 Error Detection
Vision-Language Models (VLMs) such as MiniGPT-4 100 and LLaVA 101 have demonstrated significant capabilities in understanding and interpreting images, achieving remarkable performance across various visual tasks 102, 103, 104, 105. Owing to extensive pre-training on large datasets of image-text pairs, these models possess exceptional generalization abilities, enabling impressive zero-shot performance. They can effectively recognize and process both familiar and unfamiliar objects without the need for additional training data. This capability to generalize from diverse datasets makes them highly versatile and powerful tools for a wide range of visual and multimodal applications.
@@
Our methodology employs a finely tuned prompt (SI 1) that instructs the LLM to act as a 3D printing expert agent. The VLM evaluates print quality, identifies visible failure modes, and provides structured, focused, and actionable responses in specified format.